OpenStudy (anonymous):

Evaluate the double integral of (y^-1)*(sin y)*cos(x/y) The inner integral has limits 1 and x, and has the differential dy The outer integral has limits 1 and 0, and has the differential dx I'm tearing my hair out with this one. Can anyone solve it? O.o

OpenStudy (abb0t):

\[\int\limits_{0}^{1} \int\limits_{1}^{x} (\frac{ 1 }{ y }(\sin(y)\cos(\frac{ x }{ y }))dydx \] You take the integral of the function reating x as a constant. and evanluating on 1<R<x